y=15+35.5x and y=60+28.5x
Answer by ikleyn(52781)

.
y=15+35.5x and y=60+28.5x
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

         ------------>

    15 + 35.5x = 60 + 28.5x

     35.5x - 28.5x = 60 - 15

           7x      =    45

            x      =     45/7.    ANSWER